Operations Report – February 4th

Crew 221 Operations Report 04-02-2022

SOL: 2

Name of person filing report: Luis D. Monge

Non-nominal systems: MDRS Generator is non-operational and still on inspection

Loft water tank connection presents a minor leak

Notes on non-nominal systems: Since Jan 21st, we have a 15 kW rental generator connected to our power system, meanwhile, we are working to troubleshoot the MDRS generator.

A male to male connection was bought in order to replace it in the line, but the leak it’s not in the connection, it’s in the pipe itself.

Generator: off as we are still waiting for it to reach 80%SOC to be turned on

Hours run: 12:00

From what time last night: 8:00pm

To what time this morning: 8:00am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before generator is run at night) 89%

Diesel Reading – 22%

Station Propane Reading – 65%

Ethanol Free Gasoline – gallons

Water (loft tank): 32 gallons

Water Meter: 148 units

Water (static tank) – 420 gals

Static to Loft Pump used – yes

Water in GreenHab – 267.4 gals

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: no

Hours: (before EVA) : 128.7

Beginning charge: (Before EVA)

Ending charge: (On return from EVA, before recharging)

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: no

Hours: 134,8

Beginning charge:

Ending charge:

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: Notification received on Router replacement, and it will be performed tomorrow as per instructions.

Summary of suits and radios: Radios were used during the EVAs and batteries are malfunctioning and need to be replaced.

Summary of Hab operations: during the afternoon due to high winds and low temperatures the hab temperature dropped from 22C to 18C, the wall heater in the lower deck was turned on and the temperature is back to normal, reading is 20.6C in the thermometer in the wall.

Summary of GreenHab operations: operational, watering of the plants carried out as per procedure

Summary of ScienceDome operations: 2 incubators running

Summary of RAM operations: not used

Summary of any observatory issues: no issues

Summary of health and safety issues: no issues

Questions, concerns and requests to Mission Support: a measuring cylinder and two mason jars will be needed for the science experiments, permission to use them is requested.

EVA Report – February 4th

Crew 221 EVA Report 04-02-2020

EVA # 1

Author: Jessica Todd (GreenHab Officer)

Purpose of EVA: The primary purpose was to set up several weather experiments. We were affixing two thermometers, a barometer and a rain gauge to the exterior of the Science Dome tunnel. We were also took several photos for outreach.

Start time: 12:06 pm

End time: 12:50 pm

Narrative: Today was our first maintenance EVA. The GreenHab Officer and Crew Astronomer were setting up several weather experiments as part of our goal to incorporate a Martian weather station into our MDRS mission. We spent the morning constructing a barometer and rain gauge, with the goal of affixing these and two thermometers to the exterior of the MDRS to measure weather conditions at our Hab site. We first assessed the GreenHab tunnel as a potential site for the weather experiments, however the entire tunnel is covered by a tarp and therefore rain runoff would confound our measures. We then headed to the Science Dome and found a strip of tunnel close to the Dome without tarp. We attempted to affix the thermometers to the tunnel exterior with cable ties, however the cable ties kept snapping, potentially because of the temperature, or because they have fatigued with age. We managed to affix the barometer and rain gauge to the base of the tunnel, resting on the floor, using pliers to assist in tightening the cable ties, a task we found particular difficult in our spacesuit gloves. We returned to the Hab with the thermometers having
unfortunately not met all of our objectives. We will need to determine a new method for affixing the experiments and a subsequent maintenance EVA will be required to mount the instruments.

Destination: Vicinity of the Hab.

Coordinates (use UTM NAD27 CONUS): 518500/4250050

Participants: Jessica Todd (GreenHab Officer), Ghanim Aloitabi (Astronomer)

Road(s) and routes per MDRS Map: EVA conducted near the MDRS, at the exterior of the Science Dome and GreenHab.

Mode of travel: Walking

Greenhab Report – February 4th

Crew 221 GreenHab Report 04-02-2020

GreenHab Officers: Jessica Todd

Environmental control: Heating. GreenHab door shade cloth (40%) (30%) on.

Average Temperature: 20.25°C

10:27 am

Floor Unit: 20°C

12:30 pm

Floor Unit: 22°C

1:42 pm

Floor Unit: 22°C

7:02 pm

Floor Unit: 17°C

Hours of Supplemental Light: 19:30 to 00:00

Daily water usage for crops: 17.9 gallons

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank: 267.4 gallons

Time(s) of watering for crops: 3

10:27 am = 7.9 galleons

1:42 pm = 0.5 galleons

7:05 pm = 9.5 gallons

Changes to crops: NA.

Narrative: A morning water before our maintenance EVA today. I watered a few of the herbs again in the middle of the day per Crew 220’s instructions. When watering tonight I noticed quite a few plants were dry, so perhaps too hot in the GreenHab today. We might need to open the door tomorrow for a bit to keep the temperature even. We harvested some cilantro to have with our green curry for dinner tonight.

Harvest: 3g Cilantro

Support/supplies needed: N/A

Other Notes:

Aloe and succulents unchanged. Will water again tomorrow. The strawberry plant leaves are browning and rotting at the edges. From the information sent through in yesterday’s comms window, this may be an indication of black root rot or verticillium wilt. I think I will need to check the roots tomorrow for indication of black root rot. Am not sure how to proceed in rescuing these plants. Also noticed a number of small insects around the herbs, carrots and radishes. Should I spray with pesticide?

Sol Summary – February 4th

Sol:2

Summary Title: Science Brief and a Martian friend appeared

Author’s name: Maria Grulich, M.Sc.

Mission Status: Nominal

Sol Activity Summary: Crew 221 is feeling well and woke up to a beautiful day on Mars. The weather today was windless and mostly sunny; perfect condition to set-up parts of our weather station near the hab. The whole crew was working on building our own barometer and rain gage before the crew Greenhab officer and the crew Astronomer could leave the station to set them up on the outside of the tunnels. They experienced some issues attaching the barometer as the cable ties do not seem to work great with lower temperatures on Mars. The EVA team also revealed that we have a visitor around the hab! “Fuzz” is our adorable Martian and the crew 221 made it their crew mascot which we will use for outreach activities for kids.

The whole day was under the umbrella of preparing for our science experiments. Everyone was presenting their ideas for microbiology, geology, astronomy, human factors and technology demonstrations. Luis the Crew Engineer was the most prepared with a 9-slides presentation for his projects for the crew 221. A highlight was a cinnamon apple pie prepared for the crew by the commander. Finally, the crew also decided on their call signs:

Maria -Commander will be “Ewok”

Jessica -Greenhab Officer will be “Gobblin”

Luis -Crew Engineer will be “Frankie”

Ghanim -Astronomer will be “Astro”

Hemani -Journalist Officer will be “Juno”

Rawan -HSO will be “Luna”

Tomorrow will be a busy day getting samples for microbiology and plant experiments if the EVAs will be approved. As we are coming from many different places and the crew still has to learn to understand and respect their individualities and differences, the crew 221 planned cultural nights for all 5 different countries and today will be Kuwait! We are excited about it.

Look Ahead Plan: Cultural night Kuwait, Meditation/Sport Session, EVA for microbiology and plant experiments

Anomalies in work: None

Weather: sunny no winds

Crew Physical Status: All crew in good health.

EVA: Morning 12:06-12:50 PM

Reports to be filed:

Sol Summary

Journalist Report

Photo Report

EVA Report

Operations Report

GreenHab Report

EVA Request (AM/PM)

Support Requested:

· Food request for:

· salt (empty in new food supply)

· onions (empty missing in new food supply)

· new sponges

Journalist Report – February 4th

Crew 221 Journalist Report 04Feb2020
Author: Hemani Kalucha, Crew Journalist
Sol 2
Title: Setting up a Weather Station

Today, we started with one of our science experiments: building an in-house crew assembled weather station! It actually ended up being a terrific team building exercise. We all gathered in the science dome early at 10 this morning, to put together 3 elements: a barometer, a rain gauge, and a hygrometer. Everyone put their best DIY skills to use as all of the elements were made of everyday objects found at home. Some of the intricate steps involved straws, sticky tack, cotton balls and lots of masking tape. Oh, and good timing. After this amazing team effort, we sent out two crew members on an EVA to install the 3 elements outside the hab, so that we could take weather measurements twice a day. We were pre approved to attach our instruments on the outside face of the tunnels at the science dome. Unfortunately, using zipties as a fastener was not as successful as we hoped. The EVA crew’s attempts resulted in snapping of most of the zipties. However, we have now found Velcro straps in our equipment, and will try reinstalling the equipment on a later EVA. As HabCom for this EVA, I particularly enjoyed watching the EVA crew at work from our kitchen window, and giving them special instructions for installation. Overall, I would say we had a real “starting a human presence on Mars” experience today. Unfortunately, one piece of equipment did not make it to the hab, so we will have to do without our anemometer’s wind measurements. Fortunately, Commander Grulich’s impeccable baking making skills will distract us. Dinner was made more eventful by Kuwaiti cultural night and a lovely play of “hometowns” to get to know our teammates better. Our conversations are already so much deeper and more interesting than just 48 hours ago!

Sol Summary – February 4th

Sol:2

Summary Title: Science Brief and a Martian friend appeared

Author’s name: Maria Grulich, M.Sc.

Mission Status: Nominal

Sol Activity Summary: Crew 221 is feeling well and woke up to a beautiful day on Mars. The weather today was windless and mostly sunny; perfect condition to set-up parts of our weather station near the hab. The whole crew was working on building our own barometer and rain gage before the crew Greenhab officer and the crew Astronomer could leave the station to set them up on the outside of the tunnels. They experienced some issues attaching the barometer as the cable ties do not seem to work great with lower temperatures on Mars.

The whole day was under the umbrella of preparing for our science experiments. Everyone was presenting their ideas for microbiology, geology, astronomy, human factors and technology demonstrations. Luis the Crew Engineer was the most prepared with a 9-slides presentation for his projects for the crew 221. A highlight was a cinnamon apple pie prepared for the crew by the commander. Finally, the crew also decided on their call signs:

Maria -Commander will be “Ewok”

Jessica -Greenhab Officer will be “Gobblin”

Luis -Crew Engineer will be “Frankie”

Ghanim -Astronomer will be “Astro”

Hemani -Journalist Officer will be “Juno”

Rawan -HSO will be “Luna”

Tomorrow will be a busy day getting samples for microbiology and plant experiments if the EVAs will be approved. As we are coming from many different places and the crew still has to learn to understand and respect their individualities and differences, the crew 221 planned cultural nights for all 5 different countries and today will be Kuwait! We are excited about it.

Look Ahead Plan: Cultural night Kuwait, Meditation/Sport Session, EVA for microbiology and plant experiments

Anomalies in work: None

Weather: sunny no winds

Crew Physical Status: All crew in good health.

EVA: Morning 12:06-12:50 PM

Reports to be filed:

Sol Summary

Journalist Report

Photo Report

EVA Report

Operations Report

GreenHab Report

EVA Request (AM/PM)

Support Requested:

· Food request for:

· salt (empty in new food supply)

· onions (empty missing in new food supply)

· new sponges